Given that A = 2i + 3j - k, B = i - j +2k and C = 3i + 4j + k


a)find the unit vector along the direction of nA, for n<0


b) find the length of the projection of vector A on vector C


c)find the length of projection of vector C and vector A


d) find the sine of the angle between vector A and vector C

1
Expert's answer
2022-04-12T13:52:22-0400

Given that "\overrightarrow{A} = 2\overrightarrow{i} + 3\overrightarrow{j} - \overrightarrow{k}" , "\overrightarrow{B} = \overrightarrow{i} -\overrightarrow{j} +2 \overrightarrow{k}" and "\overrightarrow{C} = 3\overrightarrow{i} + 4\overrightarrow{j}+\overrightarrow{k}" .

a)find the unit vector along the direction of nA, for n<0

"\overrightarrow{e}=-\frac{\overrightarrow{A}}{|\overrightarrow{A}|}=-\frac{2\overrightarrow{i} + 3\overrightarrow{j} - \overrightarrow{k}}{\sqrt{2^2+3^2+(-1)^2}}=""-\frac{2}{\sqrt{14}}\overrightarrow{i} -\frac{3}{\sqrt{14}}\overrightarrow{j} +\frac{1}{\sqrt{14}} \overrightarrow{k}"

b) find the length of the projection of vector A on vector C

"proj_{\overrightarrow{C}}\overrightarrow{A}=\frac{\overrightarrow{A}\cdot\overrightarrow{C}}{|\overrightarrow{C}|}=""\frac{2\cdot3+3\cdot4+(-1)\cdot1}{\sqrt{3^2+4^2+1^2}}=\frac{17}{\sqrt{26}}"

c)find the length of projection of vector C and vector A

"proj_{\overrightarrow{A}}\overrightarrow{C}=\frac{\overrightarrow{A}\cdot\overrightarrow{C}}{|\overrightarrow{A}|}=""\frac{2\cdot3+3\cdot4+(-1)\cdot1}{\sqrt{2^2+3^2+(-1)^2}}=\frac{17}{\sqrt{14}}"

d) find the sine of the angle between vector A and vector C

"|\overrightarrow{A}\times\overrightarrow{C}|=|\overrightarrow{A}|\cdot|\overrightarrow{C}|\cdot\sin{\alpha}" , where "\alpha" is the angle between "\overrightarrow{A}" and "\overrightarrow{C}"

"\sin{\alpha}=\frac{|\overrightarrow{A}\times\overrightarrow{C}|}{|\overrightarrow{A}|\cdot|\overrightarrow{C}|}"

"\overrightarrow{A}\times\overrightarrow{C}=\begin{vmatrix}
 \overrightarrow{i} & \overrightarrow{j} & \overrightarrow{k} \\
 2 & 3 & -1 \\
 3 & 4 & 1
\end{vmatrix}=""\overrightarrow{i}(3\cdot1-4\cdot(-1))-\overrightarrow{j}(2\cdot1-3\cdot(-1))+""\overrightarrow{k}(2\cdot4-3\cdot3)=""7\overrightarrow{i}-5\overrightarrow{j}-\overrightarrow{k}"

"\sin{\alpha}=\frac{\sqrt{7^2+(-5)^2+(-1)^2}}{\sqrt{2^2+3^2+(-1)^2}\cdot\sqrt{3^2+4^2+1^2}}=\frac{\sqrt{75}}{\sqrt{14}\cdot\sqrt{26}}=\frac{\sqrt{75}}{\sqrt{364}}=""\frac 52\sqrt{\frac{3}{91}}"